the perimeter of a rectangle must not be greater than 30 cm and the length must be 8cm. What is the set of values of the width?

Answer by alicealc(293)

let the length of the rectangle = l
and the width of the rectangle = w

perimeter <= 30
2*l + 2*w <= 30
(divide both sides by 2)
l + w <= 15
so, if l = 8 then
8 + w <= 15
w <= 15 - 8
w <= 7

so the width must not greater than 7 cm, or it can be said that the width is greater than 0 cm but less or equal with 7 cm (0 < width <= 7cm)